the length of a rectangle is 9cm longer then the width. The area of the rectangle is 90cm. What are the dimensions of the rectangle ?
Answer by unlockmath(1688)

Hello,
Let x be the width and x+9 be the length. The equation will be:
x(x+9)=90 Rewritten as:
x^2+9x-90=0 Factor this to:
(x-6)(x+15)=0 Solve for x:
x=6
x=-15
The obvious answer is 6 cm for the width and 15 cm for the length.
Make sense?
RJ
Check out a book I wrote at:
www.math-unlock.com